A new device ranges from $42,350-$56,210.  Your team wants the $56,210.00 version. Your facility currently bills $350 to Medicaid for 3D ultrasounds that are medically necessary and $150 for 2D ultrasounds. Private insurance is billed $385 (3D) and $185 (2D).

Analyze how many patients your facility would need to see to pay off the device in 6 months (25% of your practice is Medicaid patients) and your practice includes 5 physicians and 2 nurse-midwives. Each sees 50 patients per week.


1
Expert's answer
2021-02-01T12:24:10-0500

in six months 26 weeks

let 7 people (5+2) serve your practice 

let's make the equation

"((385-350)\times0.25\times7+(185-150)\times0.75)\times x\times26=56210"

"(35\times0.25\times7+35\times0.75)\times x\times26=56210"

"(61.25+26.25)\times x\times26=56210"

"2275\times x=56210"

x=24.7=25